OpenStudy (anonymous):

Evaluate the series: Enter a numerical answer only. ∑8k = 4(2k−1)

OpenStudy (anonymous):

Something is wrong in the question, perhaps the series is dual to 4(2n-1) such that n is the last term in the series

OpenStudy (anonymous):

Equal**

OpenStudy (anonymous):

\[\frac{ 8k \left( 8k+1 \right) }{ 2 }=4(2k-1)\]

OpenStudy (anonymous):